OpenStudy (anonymous):

evaluate the integral of e^(3x)cos(4x)dx

OpenStudy (anonymous):

integration by parts where u=cos(4x) and dv=e^3x

OpenStudy (anonymous):

\[\int\limits_{}^{}e ^{3x}(e ^{i4x}+e ^{-i4x})dx/2\]\[e ^{(3+4i)x}/(3+4i)+e ^{(3-4i)}/(3-4i)\]

OpenStudy (anonymous):

or Euler's Identity which ever one comes easier to you..LOL

OpenStudy (anonymous):

the expression will be divided by 2.

OpenStudy (anonymous):

now you can just solve to get a ans.e^3x(3cos4x+4sin3c)/25+c where c is the integration constant.

OpenStudy (anonymous):

\[\frac{1}{25} e^{3 x} (3 \text{Cos}[4 x]+4 \text{Sin}[4 x]) \]from Mathematica 8

OpenStudy (anonymous):

wow, u are so smart, you can copy answers out of a software package :/

OpenStudy (anonymous):

thank u :) u forgot the plus C robtobey

OpenStudy (anonymous):

to do it algebraically involves two steps of integration by parts

OpenStudy (anonymous):

I think they most likely wanted it done algebraically

OpenStudy (anonymous):

first video does it algebraically ( which is most likely the way they want you to do it